OpenStudy (anonymous):

A line passes through the points (-10,-4)and (-1,2) . What is the slope of the line that passes through the two points?

OpenStudy (anonymous):

\[2-(-4)/(-1)-(-10)=6/9=2/3\] the slope is 2/3

OpenStudy (anonymous):

thank you I will have to rework my problem, I did not get that

OpenStudy (anonymous):

 thats how you can find slope if you have 2 points